Find the molarity of a solution prepared by dissolving 50gram of NaOH in 100ml of water

1
Expert's answer
2021-12-01T06:26:01-0500

Molar mass of NaOH=39.997

=50/39.997= 1.25moles

= 1.25/0.1

= 12.5mol/L